# A New cw-NMR Q-meter for Dynamically Polarized Targets for Particle Physics

### v1

@@ Abstract
Polarized solid targets produced via Dynamic Nuclear Polarization rely on Continuous-Wave Nuclear Magnetism Resonance measurements to accurately determine the degree of polarization of bulk samples polarized to nearly 100%. Since the late 1970’s phase sensitive detection methods have been utilized to observe the magnetization of a sample as a small change in inductance under RF excitation near the Larmor frequency of the nuclear species of interest, using a device known as a Q-meter. Liverpool Q-meters, produced in the UK in the 80’s and 90’s, have been the workhorse devices for these targets for decades, however their age and scarcity has meant new systems are needed. We describe a Q-meter system designed and built at Jefferson Lab in the Liverpool style to have comparable electronic performance with several improvements to update and adapt the devices for modern use.
journal: Journal of Magnetic Resonance
@@ 1 Introduction
Pulsed-NMR techniques, aided by computers powerful enough to perform fast Fourier analysis, have become enormously important and broadly-applied in chemical and biological sciences [6]. However, while pulsed-NMR has the great advantage of simultaneously probing a range of frequencies to reveal the composition and structure of a material, continuous-wave NMR (cw-NMR) remains the preferred method for polarized targets, gauging the absolute degree of polarization of a given species in a bulk sample where circuit linearity limits the error achieved in a measurement. The fixed microwave pumping frequency used to drive DNP precludes the scanning of the magnetic field, necessitating instead the scanning of the RF frequency of the sampling circuit. In this article, we describe new cw-NMR instrumentation and software developed for polarized target experiments at the Thomas Jefferson National Accelerator Facility (Jefferson Lab, or JLab).
We have developed new instrumentation meant to replace the Liverpool Q-meter, the RF sampling circuit and phase sensitive millivoltmeter which form the core of a cw-NMR system. The Liverpool Q-meter has been the de facto standard for NMR in polarized targets since the 1980s, but is no longer in production and utilizes obsolete components, making maintenance difficult. The new Q-meter system described here follows the original Liverpool design closely, but takes advantage of modern components arranged in a more modular design that is intended to both reduce electromagnetic interference and facilitate easier modification and upgrade. Our system includes a new FPGA-based control and data-acquisition system for the Q-meter that permits communication over ethernet.
@@ 2 Q-meters for cw-NMR
[…]
@@ 2.7 The Construction of the Liverpool RF Module
As seen in Figure 4, the Liverpool RF module was housed in a single, machined chassis of gold-plated brass, divided into five, 2-sided, printed circuit boards with shielding partitions between each: three handling high frequency RF signals, and two low frequency (Fig. 1). The first board contains the tunable capacitance of the tank and stages of RF attenuation and amplification to beat down noise while enhancing the signal size. The second accommodates phase length tuning and holds the RF mixer, which puts out a differential low frequency signal. This differential signal is amplified on board 3. Board 4 holds the diode bridge, which also puts out a differential signal then amplified on-board. Board 5 has a final stage of low-frequency amplification for the mixer and diode signals.
@@
The chassis has penetrations for four SMA connectors, and one D-subminiature (Fig. 4). The D-sub carries the differential, low-frequency outputs for the phase and diode signals, as well as power inputs at +24, +15, $\pm$15, and +5 V. The RF signal input is supplied by an external analog signal generator, typically a Rohde & Schwarz SMT model or similar, and the input is brought into the chassis by a shielded semi-rigid cable through a separate, shielded channel directly to the mixer board, to minimize cross-talk from this largest RF signal in the device.
@@ 3.1.1 Simulation
To better understand the possible tune configurations that come with greatly increased tuning flexibility, we simulated the electronic response of a simplified Q-meter system in Python. This simulation was based initially on Mathcad code by Houlden and Court [24], and was expanded to include transmission line response from a cold tank scheme. Figure 5 shows the electronic layout of the circuit described by the simulation. The code takes the cable lengths, capacitance and phase tuning as inputs, and calculates the circuit response as both diode and phase outputs. It is available on GitHub [25] as an interactive notebook.
@@ 3.2 Q-meter Electronics
As matching the excellent performance of the Liverpool Q-meter was a chief goal, our electronic design hewed closely to the original. We emulated the steps of amplification and attenuation to keep signal levels close to those in the old boards. As in the Liverpool design, we took care to separate the high and low frequency portions of the circuit into different boards, although we split the circuit slightly differently. We used four boards: a tank board, a reference conditioning board, a high frequency “mixer” board, and a low frequency “amplifier” board. For improved shielding and modularity, each board is housed in a dedicated, EMI-shielded enclosure. These are in turn housed inside a larger aluminum chassis and interconnection between the boards are made with shielded coaxial lines.
@@ 3.2.1 High-Frequency Board
The mixer board is a two-layer board with a full ground-plane poured on each side and via-stitching connecting the two ground planes throughout. All components are surface mount, and all signal traces are on the top layer, making them effectively microstrip transmission lines. A few power supply traces jump to the bottom layer where necessary. With few exceptions, a 10 nF filter capacitor sits between every component. The 5 V DC power required by the amplifiers enters the board via 3-pin Molex nano-fit connector, and is heavily filtered by 100 $\mu$H wirewound inductors before being distributed to the amps. The ground plane is unmasked at the edges of the board make thermal and electronic contact with the chassis. Four riser bolts allow the mounting of the amplifier board on top.
@@ 3.2.3 Tank Resonator Boards
In addition to the cold tank board, a more traditional tank circuit board was available for use inside the Q-meter chassis, as seen in Figure 10. The board used the same varactors, and was connectorized with SMAs for ease of use. The board was housed in a small shielded enclosure.
@@ 3.4 Construction
To ease the adoption of our design by other groups, we aimed to use only off-the-shelf components. RF-shield enclosures from Hammond Mfg. were used, as they provided with conductive end-plate gaskets and easily accommodated Eurocard-style PCBs which slide into extruded slots. The end-plates of these enclosures were machined to accommodate SMA connectors on the front, and M12 power connectors on the back. One enclosure held the mixer and amplifier board assembly, one the reference conditioning board, and a third for the optional warm tank circuit. As seen in Figure 12, these enclosures were mounted on a plate within a 19 inch rack mount chassis, the MultipacPro from nVent Schroff, which included an EMI shielding kit to improve isolation. The power supplies were held within the same chassis, but separated from the Q-meter enclosures below the aluminium mounting plate.
@@ 4 Data Acquisition
Thr Liverpool Q-meter was often used with Yale cards to provide variable amplification, filtering and voltage offset removal, however we did not require such a system during Run Group C. This is primarily thanks to improvements in ADC resolution; our 24-bit ADCs provide sufficient accuracy at typical output voltages to obviate the need for further amplification. While we have not developed a bespoke device to amplify or remove DC voltage offsets, because our system is connectorized and modular, an SR560 low-noise preamplifier from Stanford Research Systems can be utilized. This pre-amp has capabilities beyond the Yale card and is available off-the-shelf [28].
@@ 4.2 Signal Generation
However, the Rohde & Schwarz SMA100A signal generator features an auxiliary input port, allowing faster settling times when changing the frequency. While according to the manual the device is limited to “less than” 450 $\mu$s settle time when switching via list mode, in practice we have found that the external modulation signal can be used for small hops at as fast as 50 $\mu$s without distortion in the signal. However, the SMA100A’s AUX I/O port allows direct connection to the direct digital synthesizer via a 16-bit FM data bus. This permits a very fast frequency hop with settling times of 10 $\mu$s [29]. To take advantage of the 16-bit word data bus on the SMA100A, a fourth board was built to condition 16 digital outputs from the FPGA to drive the signal generator.
@@ 4.3.1 FPGA Settings
The set registers command requires a number of settings define the frequency sweep behavior. The delay time sets how long after setting a frequency point to wait to allow the signal generator to settle. The number of samples determines how many ADC samples to read at each frequency point; this is set as a power of two to allow the averaging of the results bit-wise. The total number of sweeps to take in the sweep cycle must also be set, as well as the number of sweeps per chunk to send back intermittently. The ADC configuration register sets a number of parameters defining the capture mode and rate of the Analog Devices ADC chips.
@@ 4.4 Software
The final piece of this complete refresh of our Q-meter system was the software to coordinate the operation of the Q-meter, collect and analyze the data, and provide a user interface. Shying away from the proprietary LabVIEW software used previously, the new program was written in Python with a graphical interface based on PyQt5, which handled the full the event-driven interaction flow with threading. In addition to communicating with the FPGA data acquisition over the network, the software also directed communication and control with other crucial components of the DNP system, including microwave frequency tuning and shim magnet power supply. This software added a number of convenience and performance features, drawing on our experience with LabView NMR software used at JLab in the past. The code and documentation are available on GitHub [30].
@@
A completely new concept, the analysis tab allowed configuration of the analysis of the phase signal in three modular steps as data was being taken. The first step of the analysis requires removal of the Q-curve, the portion of the signal from the Q-meter electronics. This is typically done by taking baseline signals, in which the magnetic field is altered to shift the Larmor frequency of the spins out of the sweep frequency range. In the first modular step of the analysis, a drop-down allowed the use of a baseline or instead a polynomial fit to the background could performed and subtracted. In the second step of the analysis, any drift in the Q-curve from changes in the system was removed to completely isolate the polarization signal itself. This was done with a polynomial fit to the selectable “wings” of the signal, the portion to the left and right of the curve, excluding the polarization signal itself.
@@ 5 Experimental Performance
The first proof of the operation of our final system with polarized material came with the running of Run Group C in Hall B from June 2022 to March 2023. Two Q-meters were prepared for this experiment, each with a separate DAQ and signal generator, to allow the measurement of proton and deuteron signals at 5 T, simultaneously if desired. While the proton Q-meter channel used a tank circuit inside the Q-meter enclosure and a standard $\lambda/2$ cable, the deuteron channel took advantage of a cold tank inside the cryostat (as in Figure 9) for improved signal-to-noise with the much smaller Pake doublet signal.
@@ 5.1 Comparison with Liverpool Q-meter
To allow direct comparison of our new Q-meter system with a Liverpool Q-meter, we utilized a mechanical, single-pole double-throw RF switch, Mini-circuits RC-2SPDT-A18. This switch allowed a single tank circuit and coil to be measured by each Q-meter in turn without changing other circuit parameters. For this to be possible, it was necessary to a use a Liverpool Q-meter that was modified for an external tank circuit. In this scheme the supply and signal lines of a external tank circuit were connected to the two channels of the switch, allowing these to be connected to either setup.
@@
These tests were performed at the end of Run Group C, while the target was still operational but the experimental beam was shut down. While the time available in the hall for these tests was limited by the scheduled removal of the target system, we were able to make the necessary changes to the system to allow the comparison test. We thermalized for calibrations then polarized protons in TEMPO-doped butanol to roughly 65% positive and then negative.
@@
Figure 18 shows data from each Q-meter plotted versus the other in a range of polarizations from positive and negative. As the polarization changes quickly, particularly during the pump up or down, pairing data points from each Q-meter based on the nearest in time creates a skew away from linear. To reduce this effect in our comparison with the data taking sequence of two Liverpool points followed by two JLab points, we grouped points in four, with two points from each averaged to form a single $(x,y)$ point in this plot. So for the series of area measurements $a_{i}$ in a sequence with JLab measurements at $a_{1}$, $a_{4}$, $a_{5}$, $a_{8}$, $a_{9}$.. and Liverpool measurements at $a_{2}$, $a_{3}$, $a_{6}$, $a_{7}$, $a_{10}$.. we have:
@@ 6 Future Efforts
With the successful performance of this new Q-meter system in Hall B, we look forward to a number of improvements which will be facilitated by the modular nature of the device. A quadrature measure of both the dispersive and absorptive protions of the signal would give the complete electronic response of the resonance [34], and requires measuring both the in-phase and out-of-phase portion of the signal with an additional 90 phase shift. This would be best accomplished by replacing the diode chain of the mixer board with a second mixer; the remainder of the amplification and digitization could remain unchanged. Electronic tuning allows the synchronous tuning of the capacitance and phase shift to programmatically reduce or even remove the background Q-curve [35], and this could be implemented with updates to our VHDL code. The system’s ability to create arbitrary frequency sweep profiles is also yet to be explored. With the flexibility of using tank circuits within the cryostat, we are investigating the addition of amplification on or near the tank circuit to improve signal-to-noise.

### v2

@@ Abstract
Polarized solid targets produced via Dynamic Nuclear Polarization rely on Continuous-Wave Nuclear Magnetism Resonance measurements to accurately determine the degree of polarization of bulk samples polarized to nearly 100%. Since the late 1970’s phase sensitive detection methods have been utilized to observe the magnetization of a sample as a small change in inductance under RF excitation near the Larmor frequency of the nuclear species of interest, using a device known as a Q-meter. Liverpool Q-meters, produced in the UK in the 80’s and 90’s, have been the workhorse devices for these targets for decades, however their age and scarcity has meant new systems are needed. We describe a new cw-NMR system designed and built at Jefferson Lab, which includes a new Q-meter following the Liverpool style to have comparable electronic performance with several improvements to update and adapt the devices for modern use.
journal: Nuclear Instruments and Methods: A
@@ 1 Introduction
Pulsed-NMR techniques, aided by computers powerful enough to perform fast Fourier analysis, have become enormously important and broadly-applied in chemical and biological sciences [6]. However, while pulsed-NMR has the great advantage of simultaneously probing a range of frequencies to reveal the composition and structure of a material, continuous-wave NMR (cw-NMR) has remained the preferred method for polarized targets in scattering experiments as it provides an absorption curve that is linearly proportional to the vector polarization, without the need for Fourier transformation. Continuous-wave NMR allows the direct extraction of the polarization of quadrupolar-broadened nuclei such as 2H and 14N from the line shape [7], where in pulsed-NMR portions of the FID signal can be lost, distorting the Fourier-transformed signal [8, 9]. The fixed microwave pumping frequency used to drive DNP precludes the scanning of the magnetic field, necessitating instead scanning of the frequency of the sampling circuit.
In this article, we describe new cw-NMR instrumentation and software developed for polarized target experiments at the Thomas Jefferson National Accelerator Facility (Jefferson Lab, or JLab). This effort aimed to replace the Liverpool Q-meter, the RF sampling circuit and phase sensitive millivoltmeter which form the core of a cw-NMR system. The Liverpool Q-meter has been the de facto standard for NMR in nuclear scattering targets since the 1980s, but is no longer in production and utilizes obsolete components, making maintenance difficult. The new Q-meter system described here follows the original Liverpool design closely, but takes advantage of modern components arranged in a more modular design that is intended to both reduce electromagnetic interference and facilitate easier modification and upgrade. Our system includes a new FPGA-based control and data-acquisition system for the Q-meter that permits communication over ethernet.
@@ 2 Q-meters for cw-NMR
$P=\frac{2B_{z}}{\pi\hbar n\mu\mu_{0}}\int_{0}^{+\infty}\frac{\chi^{\prime\prime}(\omega)}{\omega}\;d\omega.$ (4)
@@ 2.7 The Construction of the Liverpool RF Module
As seen in Figure 4, the Liverpool RF module was housed in a single, machined chassis of gold-plated brass, divided into five, 2-sided, printed circuit boards with shielding partitions between each: three handling high-frequency RF signals, and two low-frequency (Fig. 1). The first board contains the tunable capacitance of the tank and stages of RF attenuation and amplification to beat down noise while enhancing the signal size. The second accommodates phase length tuning and holds the RF mixer, which puts out a differential low-frequency signal. This differential signal is amplified on board 3. Board 4 holds the diode bridge, which also puts out a differential signal then amplified on-board. Board 5 has a final stage of low-frequency amplification for the mixer and diode signals.
@@
The chassis has penetrations for four SMA connectors, and one D-subminiature (Fig. 4). The D-sub carries the differential, low-frequency outputs for the phase and diode signals, as well as power inputs at +24, +15, $\pm$15, and +5 V. The RF signal input is supplied by an external analog signal generator, typically a Rohde & Schwarz SMT model or similar, and the input is brought into the chassis by a shielded, semi-rigid cable through a separate, shielded channel directly to the mixer board to minimize cross-talk from this largest RF signal in the device.
@@ 3.1.1 Simulation
To better understand the possible tune configurations that come with greatly increased tuning flexibility, we simulated the electronic response of a simplified Q-meter system in Python. This simulation was based initially on Mathcad code by Houlden and Court [27], and was expanded to include transmission line response from a cold tank scheme. Figure 5 shows the electronic layout of the circuit described by the simulation. The code takes the cable lengths, capacitance and phase tuning as inputs, and calculates the circuit response as both diode and phase outputs. It is available on GitHub as an interactive notebook [28].
@@ 3.2 Q-meter Electronics
As matching the performance of the Liverpool Q-meter was a chief goal, our electronic design hewed closely to the original. We emulated the steps of amplification and attenuation to keep signal levels close to those in the old boards. As in the Liverpool design, we took care to separate the high and low frequency portions of the circuit into different boards, although we split the circuit slightly differently. We used four boards: a tank board, a reference conditioning board, a high frequency “mixer” board, and a low frequency “amplifier” board. For improved shielding and modularity, each board is housed in a dedicated, EMI-shielded enclosure. These are in turn housed inside a larger aluminum chassis and interconnections between the boards are made with shielded coaxial lines.
@@ 3.2.1 High-Frequency Board
[…]
@@ 3.2.3 Tank Resonator Boards
In addition to the cold tank board, a more traditional tank circuit board was available for use inside the Q-meter chassis, as seen in Figure 10. The board used the same varactors, and was connectorized with SMAs for ease of use. The board was housed in a small, shielded enclosure.
@@ 3.4 Construction
To ease the adoption of our design by other groups, we aimed to use only off-the-shelf components. RF-shield enclosures from Hammond Mfg. were used, as they were provided with conductive end-plate gaskets and easily accommodated Eurocard-style PCBs which slide into extruded slots. The end-plates of these enclosures were machined to accommodate SMA connectors on the front, and M12 power connectors on the back. One enclosure held the mixer and amplifier board assembly, one the reference conditioning board, and a third for the optional warm tank circuit. As seen in Figure 12, these enclosures were mounted on a plate within a 19 inch rack mount chassis, the MultipacPro from nVent Schroff, which included an EMI shielding kit to improve isolation. The power supplies were held within the same chassis, but separated from the Q-meter enclosures below the aluminium mounting plate.
@@ 4 Data Acquisition
The Liverpool Q-meter was often used with Yale cards to provide variable amplification, filtering and voltage offset removal, however we did not require such a system during Run Group C. This is primarily thanks to improvements in ADC resolution; our 24-bit ADCs provide sufficient accuracy at typical output voltages to obviate the need for further amplification. While we have not developed a bespoke device to amplify or remove DC voltage offsets, because our system is connectorized and modular, an SR560 low-noise preamplifier from Stanford Research Systems can be utilized. This pre-amp has capabilities beyond the Yale card and is available off-the-shelf [31].
@@ 4.2 Signal Generation
[…]
@@ 4.3.1 FPGA Settings
The set registers command requires a number of settings to define the frequency sweep behavior. The delay time sets how long after setting a frequency point to wait, allowing the signal generator to settle. The number of samples determines how many ADC samples to read at each frequency point; this is set as a power of two to allow the averaging of the results bit-wise. The total number of sweeps to take in the sweep cycle must also be set, as well as the number of sweeps per chunk to send back intermittently. The ADC configuration register sets a number of parameters defining the capture mode and rate of the Analog Devices ADC chips.
@@ 4.4 Software
The final piece of this complete refresh of our cw-NMR system was the software to coordinate the operation of the Q-meter, collect and analyze the data, and provide a user interface. Shying away from the proprietary LabVIEW software used previously, the new program was written in Python with a graphical interface based on PyQt5, which handled the full the event-driven interaction flow with threading. In addition to communicating with the FPGA data acquisition over the network, the software also directed communication and control with other crucial components of the DNP system, including microwave frequency tuning and shim magnet power supply. This software added a number of convenience and performance features, drawing on our experience with LabView NMR software used at JLab in the past. The code and documentation are available on GitHub [33].
@@
A completely new concept, the analysis tab allowed configuration of the analysis of the phase signal in three modular steps as data was being taken. The first step of the analysis requires removal of the Q-curve, the portion of the signal from the Q-meter electronics. This is typically done by taking baseline signals, in which the magnetic field is altered to shift the Larmor frequency of the spins out of the sweep frequency range. In the first modular step of the analysis, a drop-down allowed the choice of a baseline or instead a polynomial fit to the background for subtraction. In the second step of the analysis, any drift in the Q-curve from changes in the system was removed to completely isolate the polarization signal itself. This was done with a polynomial fit to the selectable “wings” of the signal, the portion to the left and right of the curve, excluding the polarization signal itself.
@@ 5 Experimental Performance
The first proof of the operation of our final system with polarized material came with the running of Run Group C in Hall B from June 2022 to March 2023. Two Q-meter systems were prepared for this experiment, each with a separate DAQ and signal generator, to allow the measurement of proton and deuteron signals at 5 T, simultaneously if desired. While the proton Q-meter channel used a tank circuit inside the Q-meter enclosure and a standard $\lambda/2$ cable, the deuteron channel took advantage of a cold tank inside the cryostat (as in Figure 9) for improved signal-to-noise with the much smaller Pake doublet signal.
@@ 5.1 Comparison with Liverpool Q-meter
To allow direct comparison of our new Q-meter system with a Liverpool Q-meter, we utilized a mechanical, single-pole double-throw RF switch, Mini-circuits RC-2SPDT-A18. This switch allowed a single tank circuit and coil to be measured by the Liverpool and JLab Q-meters in turn, without changing other circuit parameters. For this to be possible, it was necessary to a use a Liverpool Q-meter that was modified for an external tank circuit. In this scheme the supply and signal lines of a external tank circuit were connected to the two channels of the switch, allowing these to be connected to either setup.
@@
These tests were performed at the end of Run Group C, while the target was still operational but the experimental beam was shut down. While the time available in the hall for these tests was limited by the scheduled removal of the target system, we were able to make the necessary changes to the system to allow the comparison test. We thermalized for calibrations, then polarized protons in TEMPO-doped butanol to roughly 65% positive and then negative.
@@
Figure 18 shows data from each Q-meter plotted versus the other in a range of polarizations from positive and negative. As the polarization changes quickly, particularly during the pump up or down, pairing data points from each Q-meter based on the nearest in time creates a skew away from linear. To reduce this effect in our comparison, using a data taking sequence of two Liverpool points followed by two JLab points, we grouped points in four, with two points from each averaged to form a single $(x,y)$ point in this plot. So for the series of area measurements $a_{i}$ in a sequence with JLab measurements at $a_{1}$, $a_{4}$, $a_{5}$, $a_{8}$, $a_{9}$ ... and Liverpool measurements at $a_{2}$, $a_{3}$, $a_{6}$, $a_{7}$, $a_{10}$ ... we have:
@@ 6 Future Efforts
With the successful performance of this new Q-meter system in Hall B, we look forward to incorporating a number of improvements which will be facilitated by the modular nature of the device. A quadrature measure of both the dispersive and absorptive protions of the signal would give the complete electronic response of the resonance [36], and requires measuring both the in-phase and out-of-phase portion of the signal with an additional 90 phase shift. This would be best accomplished by replacing the diode chain of the mixer board with a second mixer; the remainder of the amplification and digitization could remain unchanged. Electronic tuning allows the synchronous tuning of the capacitance and phase shift to programmatically reduce or even remove the background Q-curve [37], and this could be implemented with updates to our VHDL code. The system’s ability to create arbitrary frequency sweep profiles is also yet to be explored. With the flexibility of using tank circuits within the cryostat, we are investigating the addition of amplification on or near the tank circuit to improve signal-to-noise.